# Spectral study of the pseudogap in unitary Fermi gases

### 06 Mar 2026 v1

@@ Abstract
where $\xi^{*}(\mathbf{k})\equiv\mathbf{k}^{2}/2m^{*}+U-\mu$ satisfying $\xi^{*}(k_{\mu})=0$, $\Gamma_{0}$ denotes the inverse pair lifetime, and $\Gamma_{1}$ represents the $\omega$-independent single-particle scattering rate The expression of the spectral function derived from this model via Eq. (5) is used to fit the EDCs to extract $\Delta$, $\Gamma_{0}$, and $\Gamma_{1}$. The results are shown in Fig. 5(b) and Fig. 5(c), along with the experimental data for quantitative comparison. The agreement is good. In particular, we have a finite pseudogap at and above $T_{c}$, with $\Delta\approx 0.3E_{\text{F}}$ at $T_{c}$. As shown in Fig. 5(c), the theoretical data points of $\Gamma_{0}$ (blue squares) fit nicely with an exponential, thermally activated behavior, $\Gamma_{0}\propto\exp(-2\Delta_{\text{0}}/T)$ (blue curve), suggesting that the inverse pair lifetime is dictated by the virtual pair breaking and recombination process, with excitation energy $2\Delta_{0}$. The fit yields $\Delta_{0}/E_{\text{F}}=0.41$, slightly larger than the pairing gap at $T=0.65T_{\text{c}}$, the lowest temperature accessed in experiment. The single-particle scattering rate $\Gamma_{1}$ shown in the inset of Fig. 5(c) remains nearly temperature independent, insensitive to pairing. Performing Gaussian broadening on the EDCs mainly leads to an overall increase in $\Gamma_{0}$ (and $\Gamma_{1}$) even at zero $T$, which is removed by background subtraction.
Compared with the spectral result from quantum Monte Carlo simulations [16], our present work provides more precise spectral resolutions with smaller data scatter. In comparison with other $T$-matrix based calculations, we find that the $G_{0}G_{0}$ scheme [28, 22] seems to yield an overly large pseudogap at $T_{c}$, whereas the gap seems to close at Fermi level in the $GG$ scheme [29, 10, 12, 9].
@@
In summary, we have performed a spectral study of the pseudogap phenomena in unitary Fermi gases. By going beyond previous pseudogap approximation for the self energy, and incorporating the screening to the interaction from the particle-hole channel, we have arrived at a quantitative agreement with recent experiment [15] on the spectral function, and the associated pairing gap $\Delta$, inverse pair life time $\Gamma_{0}$, and single-particle scattering rate $\Gamma_{1}$, using the same data analysis procedure as in Ref. [15]. In particular, the pseudogap self energy is calculated numerically using full convolution, which captures explicitly the Hartree shift previously dropped (or hidden in chemical potential) in the pseudogap approximation. Our work provides a microscopic explanation for the experimental data, and also serves as a support for the pairing origin of the pseudogap and for our pairing fluctuation theory.

### 07 Apr 2026 v2

@@ Abstract
where $\xi^{*}(\mathbf{k})\equiv\mathbf{k}^{2}/2m^{*}+U-\mu$ satisfying $\xi^{*}(k_{\mu})=0$, $\Gamma_{0}$ denotes the inverse pair lifetime, and $\Gamma_{1}$ represents the $\omega$-independent single-particle scattering rate. The expression of the spectral function derived from this model via Eq. (5) is used to fit the EDCs to extract $\Delta$, $\Gamma_{0}$, and $\Gamma_{1}$. The results are shown in Fig. 5(b) and Fig. 5(c), along with the experimental data for quantitative comparison. The agreement is good. In particular, we have a finite pseudogap at and above $T_{c}$, with $\Delta\approx 0.3E_{\text{F}}$ at $T_{c}$. As shown in Fig. 5(c), the theoretical data points of $\Gamma_{0}$ (blue squares) fit nicely with an exponential, thermally activated behavior, $\Gamma_{0}\propto\exp(-2\Delta_{\text{0}}/T)$ (blue curve), suggesting that the inverse pair lifetime is governed by the virtual pair breaking and recombination process, with excitation energy $2\Delta_{0}$. The fit yields $\Delta_{0}/E_{\text{F}}=0.41$, slightly larger than the pairing gap at $T=0.65T_{\text{c}}$, the lowest temperature accessed in experiment. The single-particle scattering rate $\Gamma_{1}$ shown in the inset of Fig. 5(c) remains nearly temperature independent, insensitive to pairing. Performing Gaussian broadening on the EDCs mainly leads to an overall increase in $\Gamma_{0}$ (and $\Gamma_{1}$) even at zero $T$, which is removed by background subtraction.
@@
In summary, we have performed a spectral study of the pseudogap phenomena in unitary Fermi gases. By going beyond the previous pseudogap approximation for the self-energy, and incorporating the particle-hole channel screening to the interaction, we have arrived at a quantitative agreement with the recent experiment [14] on the spectral function, and the associated pairing gap $\Delta$, inverse pair lifetime $\Gamma_{0}$, and single-particle scattering rate $\Gamma_{1}$, using the same data analysis procedure as in Ref. [14]. In particular, the pseudogap self-energy is calculated numerically using full convolution, which captures explicitly the Hartree shift previously dropped (or hidden in chemical potential) in the pseudogap approximation. Our work thus provides a quantitative microscopic explanation for the observed pseudogap and establishes pairing fluctuations as the key mechanism underlying the pseudogap in unitary Fermi gases.